1.The nurse is to administer an IM injection to the patient who weighs 125 kg. The nurse is correct to choose which of the following needles for the injection? A. 23G 2" , B. 22G 1" , C. 22G 1 1/2" , D. 21G 5/8"

2. The nurse is to administer an intradermal injection to the patient. The ordered dose is contained in 0.2 mL. The nurse is correct to choose which of the following syringes and needles for the injection? (Select all that apply.) A. 1 mL 26G 3/8” , B. 3 mL 22G 5/8” , C. 1 mL 26G 1/2” , D. 3 mL 25G 5/8” , E. 1 mL 28G 1/2”

3. The nurse reconstituted a 10 g vial of powdered medication. The instructions for reconstitution state: Add 10 mL, 5 mL, or 2.5 mL of Sterile Water for Injection to provide 1 g, 500 mg, or 250 mg per mL, respectively. The nurse added 5 mL to the vial. What is the dosage strength of the medication that the nurse prepared? A.10 g / 5 mL , B. 500 mg / 5 mL , C.500 mg / mL , D. 10 g / 2.5 mL

4.

The physician orders 500 mg of cefepime IV q.12h. The nurse has a 2 g vial of cefepime with the following instructions for reconstitution:

Single-dose vials for IV/IM administration

Amount of diluent to be added (mL)

Approximate available volume (mL)

Approximate
cefepine concentration (mg / mL)

1 g (IV)

10

11.3

100

1 g (IM)

2.4

3.6

280

2 g (IV)

10

12.5

160


The nurse is correct to:

A. add 2.4 mL diluent to the powder and administer 1.79 mL to the patient.

B. add 10 mL diluent to the powder and administer 3.13 mL to the patient.

C.add 10 mL diluent to the powder and administer 2.5 mL to the patient.

D. add 10 mL diluent to the powder and administer 5 mL to the patient.
Answer #1

1. For IM injection to the patient who weighs 125 kgs the needle for injection should be 22G 1 1/2". Option C is the correct answer.

2. For an intradermal injection to the patient with 0.2 ml dose 1 mL 26G 1/2” should be used. Option C is the correct answer.

3. The nurse reconstituted a 10 g vial with 5 ml of sterile water. The dosage strength of the medication that the nurse prepared is 10 g / 5 mL. Option A is the correct answer.

4. Prescribed 500 mg of cefepime IV q.12h.

Medicine available on hand is 2 g vial of cefepime

The volume of medicine needed = required dose \times available volume \div available dose

                                             = 500 mg \times 12.5 ml \div 2000 mg (2gms)

                                             = 3.125 (can be rounded to 3.13 ml)

The nurse should add 10 mL diluent to the powder and administer 3.13 mL to the patient. Option B is the correct answer.